# Exact and Linear Convergence for Federated Learning under Arbitrary Client Participation is Attainable

### v2

@@ Abstract
This work tackles the fundamental challenges in Federated Learning (FL) posed by arbitrary client participation and data heterogeneity, prevalent characteristics in practical FL settings. It is well-established that popular FedAvg-style algorithms struggle with exact convergence and can suffer from slow convergence rates since a decaying learning rate is required to mitigate these scenarios. To address these issues, we introduce the concept of stochastic matrix and the corresponding time-varying graphs as a novel modeling tool to accurately capture the dynamics of arbitrary client participation and the local update procedure. Leveraging this approach, we offer a fresh perspective on designing FL algorithms, provide a rigorous quantitative analysis of the limitations inherent in the FedAvg algorithm, and present FOCUS , Federated Optimization with Exact Convergence via Push-pull Strategy, a provably convergent algorithm designed to effectively overcome the previously mentioned two challenges. More specifically, we provide a rigorous proof demonstrating that FOCUS achieves exact convergence with a linear rate regardless of the arbitrary client participation, establishing it as the first work to demonstrate this significant result.
@@ 1 Introduction
Federated Learning has emerged as a powerful paradigm for distributed machine learning, enabling multiple clients to collaboratively train models without sharing raw data. However, a central challenge in FL is the arbitrary and unpredictable nature of client participation. In real-world FL, clients may join or leave at will, participate intermittently, or drop out due to connectivity or resource constraints.
@@
where $x\in{\mathbb{R}}^{d}$ represents the $d-$dimensional model parameter and $f_{n}$ stands for the local cost function. It is well established that when clients perform multiple local updates on non-i.i.d. data, their local models tend to diverge. This leads to client drift from the optimal solution of problem (1), a phenomenon that persists even under the often impractical uniform client sampling assumption (Karimireddy et al., 2020; Li et al., 2019a). Moreover, arbitrary client participation introduces another objective bias: instead of converging to the true global optimum, the global model converges to a stationary point of a distorted, participation-weighted objective (Wang and Ji, 2022; Wang et al., 2020). To mitigate this persistent error, existing methods typically require decaying the learning rate asymptotically to zero, at least in theory. While this strategy can reduce the bias in the limit, it often leads to slower convergence. Hence, a key question naturally arises:
@@
We will provide an affirmative answer to this question in this paper. We begin by introducing a novel analytical framework that reformulates the core operations of FL—client participation, local updates, and model aggregation over time-varying graphs—as a sequence of stochastic matrix multiplications (Horn and Johnson, 2012). Next, with this tool, we develop a new algorithm FOCUS , Federated Optimization with Exact Convergence via Push-pull Strategy, which is inspired by decentralized optimization algorithms (Nedic and Ozdaglar, 2009; Sayed et al., 2014; Lian et al., 2017; Lan et al., 2020). More specifically, we leverage the push-pull technique (Pu et al., 2020; Xin and Khan, 2018) with the time-varying graphs (Nedic et al., 2017; Ying et al., 2021; Nguyen et al., 2025) instead of commonly used static or strongly connected communication graphs, since FOCUS is designed for the FL setting. Compared to the variance reduction technique (Johnson and Zhang, 2013; Defazio et al., 2014) or the adaptively learning participation probabilities, the push-pull approach handles the unknown client participation scenario much better both empirically and theoretically.
@@
  • Even under arbitrary client participation, FOCUS exhibits linear convergence (exponential decay) for both convex and non-convex (with PL condition) scenarios without assuming the bounded heterogeneity or decaying the learning rates.
@@
Algorithm Exact Converg.1 Strongly-Convex Complexity2 Non-Convex Complexity Assumptions5
@@
FedAvg (Li et al., 2019a) $O(\frac{1}{\epsilon})$ $O(\frac{1}{\epsilon^{2}})$ Uniform Bounded Bounded gradient assumption
LocalSGD (Koloskova et al., 2020) $O(\frac{1}{\sqrt{\epsilon}})$ $O(\frac{1}{\epsilon^{3/2}})$ Uniform Bounded Doubly stochastic matrix
FedAU (Wang and Ji, 2023) $O\big{(}\frac{1}{\epsilon^{2}}\big{)}$ Arbitrary Bounded Bounded global gradient
FedAWE (Xiang et al., 2024) $O(\frac{1}{\epsilon^{2}})$ Arbitrary Bounded Doubly stochastic matrix
@@
ProxSkip/ScaffNew Mishchenko et al. (2022) $O\big{(}\log(\frac{1}{\epsilon})\big{)}$ Full None Comm. $2d$ vector per round
FOCUS (This paper) $O\big{(}\log(\frac{1}{\epsilon})\big{)}$ $O\big{(}\log(\frac{1}{\epsilon})\big{)}^{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4}$ Arbitrary None No need to learn partici. prob.
Table 1: Comparison of multiple algorithms. 1 Exact convergence refers to the algorithm’s ability to converge to the exact solution under arbitrary sampling, without requiring a decaying learning rate. 2 Complexity refers to the number of iterations required for the algorithm to achieve an error within $\epsilon$ of the optimal solution. We have removed the impact of the stochastic gradient variance in all rates. 3 There is no convergence proof of SCAFFOLD under arbitrary client participation scenario. Empirically, we observed it may be possible. 4 This rate is established with PL condition. 5 Arbitrary participation refers to Assumption 1 and the bounded heterogeneous gradient are the assumptiopns that $\|f_{i}(x)-F(x)\|\leq\sigma_{G}$.
@@ 2 Related Work
For $\mbox{{For}\; }t=0,1,\cdots,\tau-1:$ (Local Update)
@@
$x_{r+1}\Leftarrow\frac{1}{|S_{r}|}\sum_{i\in S_{r}}x_{\tau+1,i}^{(r)}$ (Aggregate Model) (2c)
@@
Because of the data heterogeneity and multiple local update steps, Li et al. (2019a) has shown that the fixed point of FedAvg is not the same as the minimizer of (1) in the convex scenario. More specifically, they quantified that
$\|x^{o}-x^{\star}\|^{2}=\Omega\big{(}(\tau-1)\eta\big{)}\|x^{\star}\|^{2},$ (3)
@@
Many analytical studies on FL assume that the sampled clients is drawn from a uniform distribution, an assumption shared by the literature cited in the preceding paragraph, but this is almost impractical in reality (Kairouz et al., 2021; Xiang et al., 2024). Wang and Ji (2022) demonstrated that FedAvg might fail to converge to $x^{\star}$ under non-uniform sampling distributions, even with a decreasing learning rate $\eta$. To address the challenges posed by non-uniformity, a common approach involves either explicitly knowing or adaptively learning the client participation probabilities during the iterative process and subsequently modifying the averaging weights accordingly (Wang and Ji, 2023; Wang et al., 2024; Xiang et al., 2024). However, neither of them can achieve exact convergence, and the learning process may slow down the convergence.
It is known that federated learning (FL) and decentralized optimization algorithms are closely related (Lalitha et al., 2018; Kairouz et al., 2021; Koloskova et al., 2020). This work is closely related to the tools introduced in the decentralized optimization society. We leave a more detailed decentralized literature review in the Appendix A.
@@ 3 Graph, Stochastic Matrix, and Arbitrary Client Participation
$R(S_{r})[i,j]=\begin{cases}1&\mbox{if $i\in S_{r}$ and $j=0$}\\ 1&\mbox{if $i\notin S_{r}$ and $j=i$}\\ 0&\mbox{otherswise}\end{cases},\;\;\;A(S_{r})[i,j]=$$\begin{cases}1&\mbox{if $i=j\neq 0$}\\ 1/|S_{r}|&\mbox{if $i\in S_{r}$ and $j=0$}\\ 0&\mbox{otherswise}\end{cases}$ (4)
@@
$C(S_{r})[i,j]\!=\!\begin{cases}1&\mbox{if $j\in S_{r}$ and $i=0$}\\ 1&\mbox{if $j\notin S_{r}$ and $i=j$}\\ 0&\mbox{otherswise}\end{cases},\;\;\;W(S_{r})[i,j]\!=\!\begin{cases}1/|S_{r}|&\mbox{if $i\in S_{r}$ and $j=0$}\\ 1/|S_{r}|&\mbox{if $j\in S_{r}$ and $i=0$}\\ 1-\sum_{i}{W[i,j]}&\mbox{if $i=j$}\\ 0&\mbox{otherswise}\end{cases}$
Suppose $S_{r}=\{1,3\}$, then the matrices $W(S_{r})$ and $C(S_{r})$ correspond to the rightmost and second rightmost ones depicted in the figure. These four matrices will play the critical role in the following algorithm design and convergence proof section. In contrast to decentralized algorithms, where assumptions are directly imposed on the mixing matrix, we do not make any assumption about them in this paper since we utilize them to model the client participation process.
@@ 3.1 Arbitrary Client Participation Modeling
The column-stochastic matrix equals ($\bar{C}=\mathbb{E}[C(S_{r})]=\bar{R}^{\mathsf{T}}$) by definition. While doubly stochastic matrices are prevalent in the literature, they are not applicable to FL-style algorithms, and therefore we do not discuss them further. With this approach, we effectively transform the problem of arbitrary client participation probabilities into an analysis of the matrix properties of $R(S_{r})$, $A(S_{r})$, and $C(S_{r})$ as which we will exploit in the subsequent section.
@@ 4.1 Interpret FedAvg as Decentralized Algorithm with Time-Varying Graphs
A direct application of the above mixing matrix is that we can concisely represent the FL algorithm in vector-matrix form, similar to decentralized algorithms (Li et al., 2019a; Koloskova et al., 2020).
First, let ${\boldsymbol{x}}_{k}={\rm vstack}[x_{k,0};x_{k,1};\cdots;x_{k,N}]\in\mathbb{R}^{(N+1)\times d}$ denote the state at iteration $k$. This matrix is formed by vertically stacking the server’s model parameters $x_{k,0}\in\mathbb{R}^{d}$ and the model parameters $x_{k,i}\in\mathbb{R}^{d}$ from the $N$ workers. Similarly, let $\nabla{\boldsymbol{f}}({\boldsymbol{x}}_{k})={\rm vstack}[\mathbb{0};\nabla f_{1}(x_{k,1});\cdots;\nabla f_{N}(x_{k,N})]\in\mathbb{R}^{(N+1)\times d}$ represent the corresponding stacked vector of local gradients at iteration $k$ [^1] In the decentralized optimization literature, it is common to represent parameters $x$ and gradients $\nabla f(x)$ as row vectors (dimension $1\times d$). This allows the graph mixing operation, defined by a matrix $W\in\mathbb{R}^{(N+1)\times(N+1)}$, to be concisely expressed as $W{\boldsymbol{x}}_{k}$. Note that the first component of the stacked gradient is $\mathbf{0}$ because the server holds no data. This implies that the server’s local loss function is identically zero, $f_{0}(x)\equiv 0$, and consequently, $\nabla f_{0}(x_{k,0})=\mathbb{0}$. This also ensures that including the server’s term $f_{0}$ does not alter the original loss function defined in (1).
@@
This diagonal matrix $D_{k}$ serves to deactivate unparticipated workers and the server during local updates. $S_{r}$ is the set of participated clients’ indices at round $r$, which can be determined by the iteration $k$, i.e., $r\tau\leq k<(r+1)\tau$. An illustration this process using graphs are shown in Figure 2.
@@
Convergence Result of FedAvg with Arbitrary Participations. In the appendix C, we provide a new proof of FedAvg under the arbitrary participation scenario through this decentralized optimization formulation. When the algorithm $k\to\infty$, the limiting point of FedAvg is around an irreducible neighborhood depending on the local update steps $\tau$, data heterogeneity $\sigma_{g}^{2}$, and the extra bias $\delta_{q}^{2}$ introduced due to non-uniform participation probabilities. This motivates us to develop a new FL algorithm capable of addressing and eliminating all aforementioned errors and biases.
@@
Figure 3: Illustration of our new
FOCUS
algorithm. There are two key differences from FedAvg style algorithm. One is it pulls the model variable ${\boldsymbol{x}}$ but pushes the gradient variable ${\boldsymbol{y}}$, and another is the push matrix is the column stochastic matrix instead of the row stochastic.
4.2 FOCUS Corrects the Arbitrary Client Participation and Local-Update Bias
4.2.1 Push-Pull strategy for FL setting

### v3

@@ Abstract
This work tackles the fundamental challenges in Federated Learning (FL) posed by arbitrary client participation and data heterogeneity, prevalent characteristics in practical FL settings. It is well-established that popular FedAvg-style algorithms struggle with exact convergence and can suffer from slow convergence rates since a decaying learning rate is required to mitigate these scenarios. To address these issues, we introduce the concept of stochastic matrix and the corresponding time-varying graphs as a novel modeling tool to accurately capture the dynamics of arbitrary client participation and the local update procedure. Leveraging this approach, we offer a fresh decentralized perspective on designing FL algorithms and present FOCUS, Federated Optimization with Exact Convergence via Push-pull Strategy, a provably convergent algorithm designed to effectively overcome the previously mentioned two challenges. More specifically, we provide a rigorous proof demonstrating that FOCUS achieves exact convergence with a linear rate regardless of the arbitrary client participation, establishing it as the first work to demonstrate this significant result.
@@ 1 Introduction
Federated Learning (FL) has emerged as a powerful paradigm for distributed learning, enabling multiple clients to collaboratively train models without sharing raw data. Yet, a central challenge in FL is the arbitrary and unpredictable nature of client participation. In real-world FL, clients may join or leave at will, participate intermittently, or drop out due to connectivity or resource constraints.
@@
where $x\in{\mathbb{R}}^{d}$ represents the $d$-dimensional model parameter and $f_{n}$ stands for the local cost function. It is well established that when clients perform multiple local updates on non-i.i.d. data, their local models tend to diverge. This leads to client drift from the optimal solution of problem (1), a phenomenon that persists even under the often impractical uniform client sampling assumption (Karimireddy et al., 2020; Li et al., 2020). Moreover, arbitrary client participation introduces another objective bias: instead of converging to the true global optimum, the global model converges to a stationary point of a distorted, participation-weighted objective (Wang et al., 2020; Wang and Ji, 2022). To mitigate this persistent error, existing methods typically require decaying the learning rate asymptotically to zero, at least in theory. While this strategy can reduce the bias in the limit, it often leads to slower convergence. Hence, a key question naturally arises:
@@
We will provide an affirmative answer to this question in this paper. We begin by introducing a novel analytical framework that reformulates the core operations of FL - client participation, local updates, and model aggregation over time-varying graphs - as a sequence of stochastic matrix multiplications (Horn and Johnson, 2012). Next, with this tool, we develop a new algorithm FOCUS, Federated Optimization with Exact Convergence via Push-pull Strategy, which is inspired by decentralized optimization algorithms (Nedic and Ozdaglar, 2009; Sayed et al., 2014; Lian et al., 2017; Lan et al., 2020). More specifically, we leverage the push-pull technique (Xin and Khan, 2018; Pu et al., 2020) with the time-varying graphs (Nedic et al., 2017; Ying et al., 2021; Nguyen et al., 2025) instead of commonly used static or strongly connected communication graphs, since FOCUS is designed for the FL setting. Compared to the variance reduction technique (Johnson and Zhang, 2013; Defazio et al., 2014) or the adaptively learning participation probabilities, the push-pull approach handles the unknown client participation scenario much better both empirically and theoretically.
@@
  • Even under arbitrary client participation, FOCUS exhibits linear convergence (exponential decay) for both strongly convex and non-convex (with PL condition) scenarios without assuming the bounded heterogeneity or decaying the learning rates.
@@
Algorithm
Exact
Converg.1
Strongly-Convex
Complexity2
Non-Convex
Complexity
Assumptions5
@@
FedAvg
(Li et al., 2020)
$O(\frac{1}{\epsilon})$ $O(\frac{1}{\epsilon^{2}})$ Uniform Bounded Bounded gradient assumption
LocalSGD
(Koloskova et al., 2020)
$O(\frac{1}{\sqrt{\epsilon}})$ $O(\frac{1}{\epsilon^{3/2}})$ Uniform Bounded Doubly stochastic matrix
FedAU
(Wang and Ji, 2024)
$O\big(\frac{1}{\epsilon^{2}}\big)$ Arbitrary Bounded Bounded global gradient
FedAWE
(Xiang et al., 2024)
$O(\frac{1}{\epsilon^{2}})$ Arbitrary Bounded Doubly stochastic matrix
SCAFFOLD
(Karimireddy et al., 2020)
@@
ProxSkip/ScaffNew
(Mishchenko et al., 2022)
$O\big(\log(\frac{1}{\epsilon})\big)$ Full None Comm. $2d$ vector per round
MIFA
(Gu et al., 2021)
- $O(\frac{1}{\epsilon^{2}})$ Arbitrary Bounded Bounded delay Assump. + Server stores each client model
FOCUS
(This paper)
$O\big(\log(\frac{1}{\epsilon})\big)$ $O\big(\log(\frac{1}{\epsilon})\big)^{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4}$ Arbitrary None No need to learn partici. prob.
Table 1: Comparison of multiple algorithms. 1 Exact convergence refers to the algorithm’s ability to converge to the exact solution under arbitrary sampling, without requiring a decaying learning rate. 2 Complexity refers to the number of iterations required for the algorithm to achieve an error within $\epsilon$ of the optimal solution. We have removed the impact of the stochastic gradient variance in all rates. 3 There is no convergence proof of SCAFFOLD under arbitrary client participation scenario. Empirically, we observed it may be possible. 4 This rate is established with PL condition. 5 Arbitrary participation refers to Assumption 1 and the bounded heterogeneous gradient are the assumptions that $\|f_{i}(x)-F(x)\|\leq\sigma_{G}$. 6 It is possible to reduce the uplink communication into $d$ while downlink one is still $2d$ (Huang et al., 2024).
@@ 2 Related Work
$\mbox{{For}\; }t=0,1,\cdots,\tau-1:$ (Local Update)
@@
$x_{r+1}\Leftarrow\frac{1}{|S_{r}|}\sum_{i\in S_{r}}x_{\tau,i}^{(r)}$ (Aggregate Model) (2c)
@@
Because of the data heterogeneity and multiple local update steps, Li et al. (2020) has shown that the fixed point of FedAvg is not the same as the minimizer of (1) in the convex scenario. More specifically, they quantified that
$\vskip-8.53581pt\|x^{o}-x^{\star}\|^{2}=\Omega\big((\tau-1)\eta\big)\|x^{\star}\|^{2},\vskip-8.53581pt$ (3)
@@
Many analytical studies on FL assume that the sampled clients are drawn from a uniform distribution, an assumption shared by the literature cited in the preceding paragraph, but this is almost impractical in reality (Kairouz et al., 2021; Xiang et al., 2024; Li et al., 2025). Wang and Ji (2022) shows that FedAvg might fail to converge to $x^{\star}$ under non-uniform sampling distributions, even with a decreasing learning rate $\eta$. To address the challenges posed by non-uniformity, a common approach involves either explicitly knowing or adaptively learning the client participation probabilities during the iterative process and subsequently modifying the averaging weights accordingly (Wang and Ji, 2024; Wang et al., 2024; Xiang et al., 2024). Yet, neither of them can achieve exact convergence, and the learning process may slow down the convergence. An alternative approach is to use Variance Reduction (VR) techniques, as seen in methods like MIFA Gu et al. (2021) and FedVARP Jhunjhunwala et al. (2022). Yet, the heuristic integration of VR with FL often fails to jointly address the client drift issue. This results, once again, in inexact convergence when a constant learning rate is used.
It is known that FL and decentralized optimization are closely related (Lalitha et al., 2018; Koloskova et al., 2020; Kairouz et al., 2021), and this work is closely related to the tools introduced in the decentralized optimization society. We leave a detailed decentralized literature review in Appendix A.
@@ 3 Graph, Stochastic Matrix, and Arbitrary Client Participation
$R(S_{r})[i,j]=\begin{cases}1&\mbox{if $i\in S_{r}$ and $j=0$}\\ 1&\mbox{if $i\notin S_{r}$ and $j=i$}\\ 0&\mbox{otherwise}\end{cases},\;\;\;A(S_{r})[i,j]=$ $\begin{cases}1&\mbox{if $i=j\neq 0$}\\ 1/|S_{r}|&\mbox{if $i\in S_{r}$ and $j=0$}\\ 0&\mbox{otherswise}\end{cases}$ (4)
@@
[…]
Suppose $S_{r}=\{1,3\}$, then the matrices $W(S_{r})$ and $C(S_{r})$ correspond to the rightmost and second rightmost ones depicted in the figure. These four matrices will play the critical role in the following algorithm design and convergence proof section. In contrast to decentralized algorithms, where assumptions are directly imposed on the mixing matrix, we do not make any assumption about them in this paper since we utilize them to model the client participation process. For completeness, a brief review of stochastic matrices and their properties is provided in the Appendix C.
@@ 3.1 Arbitrary Client Participation Modeling
The column-stochastic matrix equals ($\bar{C}=\mathbb{E}[C(S_{r})]=\bar{R}^{\mathsf{T}}$) by definition. While doubly stochastic matrices are prevalent in the decentralized literature, they are not often applicable to FL-style algorithms, and therefore we do not discuss them further. With this approach, we effectively transform the problem of arbitrary client participation probabilities into an analysis of the matrix properties of $R(S_{r})$, $A(S_{r})$, and $C(S_{r})$ as which we will exploit in the subsequent section.
@@
In this section, we demonstrate how leveraging the graph and stochastic matrix can facilitate the development of more powerful FL algorithms.
@@ 4.1 Interpret FedAvg as Decentralized Algorithm with Time-Varying Graphs
A direct application of the above mixing matrix is that we can concisely represent the FL algorithm in vector-matrix form, similar to decentralized algorithms (Li et al., 2020; Koloskova et al., 2020).
First, let ${\boldsymbol{x}}_{k}={\rm vstack}[x_{k,0};x_{k,1};\cdots;x_{k,N}]\in\mathbb{R}^{(N+1)\times d}$ denote the state at iteration $k$. This matrix is formed by vertically stacking the server’s model parameters $x_{k,0}\in\mathbb{R}^{d}$ and the model parameters $x_{k,i}\in\mathbb{R}^{d}$ from the $N$ workers. Similarly, let $\nabla{\boldsymbol{f}}({\boldsymbol{x}}_{k})={\rm vstack}[\mathbb{0};\nabla f_{1}(x_{k,1});\cdots;\nabla f_{N}(x_{k,N})]\in\mathbb{R}^{(N+1)\times d}$ represent the corresponding stacked vector of local gradients at iteration kk. [^1] In the decentralized optimization literature, it is common to represent parameters $x$ and gradients $\nabla f(x)$ as row vectors (dimension $1\times d$). This allows the graph mixing operation, defined by a matrix $W\in\mathbb{R}^{(N+1)\times(N+1)}$, to be concisely expressed as $W{\boldsymbol{x}}_{k}$. Note that the first component of the stacked gradient is $\mathbb{0}$ because the server holds no data. This implies that the server’s local loss function is identically zero, $f_{0}(x)\equiv 0$, and consequently, $\nabla f_{0}(x_{k,0})=\mathbb{0}$. This also ensures that including the server’s term $f_{0}$ does not alter the original loss function defined in (1).
@@
This diagonal matrix $D_{k}$ serves to deactivate unparticipated workers and the server during local updates. $S_{r}$ is the set of participated clients’ indices at round $r$, which can be determined by the iteration $k$, i.e., $r\tau\leq k<(r+1)\tau$. An illustration of this process using graphs is shown in Figure 2.
@@
Convergence Result of FedAvg with Arbitrary Participations. In the appendix D, we provide a new proof of FedAvg under the arbitrary participation scenario through this decentralized optimization formulation. When the algorithm $k\to\infty$, the limiting point of FedAvg is around an irreducible neighborhood depending on the local update steps $\tau$, data heterogeneity $\sigma_{g}^{2}$, and the extra bias $\delta_{q}^{2}$ introduced due to non-uniform participation probabilities. This motivates us to develop a new FL algorithm capable of addressing and eliminating all aforementioned errors and biases.
@@
Figure 3: Illustration of our new FOCUS algorithm. There are two key differences from FedAvg style algorithm. One is it pulls the model variable $\boldsymbol{x}$} but pushes the gradient variable $\boldsymbol{y}$}, and another is the push matrix is the column stochastic matrix instead of the row stochastic.
4.2 FOCUS Corrects Arbitrary Client Participation and Local-Update Bias
4.2.1 Push-Pull Strategy for FL Settings